# A-H Premium and the Shanghai-Hong Kong Stock Connect

### v1

@@ 3 Empirical results
We first test Hypothesis 1, which examines the effect of the SHHK Stock Connect on the A-H share price premium. Table 2 reports the baseline dynamic panel regression results. Column (1) only includes the lagged premium and the SHHK Stock Connect indicator. The coefficient on $SHHKPolicy$ is positive and statistically significant at the 1% level, indicating that the introduction of the Stock Connect is associated with an increase in the A-H share price premium.
Columns (2) to (6) progressively incorporate additional firm-level control variables, with Column (6) representing the full specification. Across all specifications, the estimated coefficient on $SHHKPolicy$ remains positive and significant at 1% level. In the full model, the coefficient on $SHHKPolicy$ is 0.184, implying that, holding other variables constant, the implementation of the SHHK Stock Connect is associated with an average increase of approximately 18.4% in the A-H share price premium for a given pair of dual-listed shares. The baseline regression results present an average effect of the SHHK Stock Connect on the A-H share price premium. However, the economic mechanisms under this effect remain unclear. If SHHK Stock Connect operates by alleviating market frictions and improving information transmission, its impact should be more pronounced in less efficient markets. This motivates our analysis of efficiency-dependent effects in the next subsection.
@@
3.1 Market Efficiency and the Effects of the SHHK Stock Connect
@@
Meng et al., (2023) finds that the SHHK Stock Connect increases the A share market efficiency of AH dual listing firms through stock price informational efficiency and corporate governance. Motivated by their research, we test whether the effect of the SHHK Stock Connect on the A-H share price premium depends on market efficiency. We construct the following model:
$\text{Prem}_{it}=\alpha_{i}+\beta_{1}\text{Prem}_{i,t-1}+\beta_{2}\text{SHHKPolicy}_{t}+\beta_{3}\text{Effboth}_{t}+\beta_{4}\text{SHHKPolicy $\times$ Effboth}_{t}+\gamma^{\prime}X_{it}+\varepsilon_{it}$ (8)
where the $Prem_{i,t}$ is the AH share price premium for share $i$ at month $t$. Follow Domowitz et al., (1997) and Fan and Wang, (2017), we control the lagged term $Prem_{i,t-1}$ to reduce the effect of autocorrelation.
Table 3 reports the regression results. Interestingly, after considering the interaction term between the Stock Connect indicator and market efficiency ($SHHKPolicy\times Effboth$) and efficiency for both market ($Effboth$), the coefficient of $SHHKPolicy$ becomes negative and statistically significant. The reason is that coefficient on $SHHKPolicy$ no longer captures the average policy effect, but rather the effect evaluated at $Effboth=0$. Since $Effboth$ is strictly negative in our sample, this coefficient should not be interpreted in isolation. In this model, the marginal effect of SHHK Stock Connect is:
$\frac{\partial prem}{\partial SHHKPolicy}=\beta_{2}+\beta_{4}Effboth,$
so that, under our scaling where $Effboth<0$ and values closer to zero indicate higher efficiency, the estimated policy effect increases as $Effboth$ becomes more negative (i.e., as efficiency declines). Consequently, the Stock Connect is associated with a larger change in the AH premium for firms operating under less efficient market conditions than for those with higher efficiency. Overall, the results in Table 3 support an efficiency-dependent policy effect both statistically and economically.
@@
Lagged Premium 0.633***
SHHKPolicy $0.490^{**}$
Effboth 39.554**
SHHKPolicy $\times$ Effboth $45.900^{***}$
Turnover 0.022***
@@
Demand 0.009-0.009
Firm Size $0.108^{***}$
Interest Rate 0.007***
@@ 3.2 Robustness check
Table 5 examines whether the announcement of the SHHK Stock Connect has an efficiency-dependent effect on the A-H price premium. The coefficient on $Announcement$ is negative but statistically insignificant (0.217-0.217), suggesting no significant change in the AH premium around the announcement date. More importantly, the interaction term AnnouncementAnnouncement $\times$ EffbothEffboth is also statistically insignificant (14.471-14.471), providing no evidence that the announcement effect varies systematically with market efficiency. Consistent with these results, the main effect of $Effboth$ is not statistically different from zero in this specification. Overall, Table 5 indicates that the announcement itself and the interaction with market efficiency are unlikely to generate a measurable impact on the AH premium.
@@
According to Meng et al., (2023), we use an alternative market efficiency proxy, $Effboth2$, as a robustness check. Specifically, for each trading day $d$ we compute the relative high-low spread as
$S_{2}=\frac{P^{high}_{d}-P^{low}_{d}}{\left(P^{high}_{d}+P^{low}_{d}\right)/2},$ (9)
@@
$Eff_{2}=-S_{2},$ (10)
For A share market and H share market, we calculate the $Eff_{2}$ separately and then add them up as an alternative proxy for overall market efficiency ($EFFboth2$). Table 6 presents our regression results for third robustness check. Even though $SHHKPolicy$ is not statistically significant, the sign of coefficient is negative. The interaction term $SHHKPolicy\times Effboth2$ is negatively significant at 5% level. So, in general, this result is consistent with the result shown in Table 3.
@@
The SHHK Stock Connect was launched in November, 2014, to further test whether the effect comes from SHHK Stock Connect and assess the robustness of our previous findings, we conduct a placebo test by assigning the implementation of the SHHK Stock Connect to a later date. Specifically, we assume that the policy was introduced two years after its actual launch and construct a binary indicator, $FakePolicy$, which equals one for observations after November 2016 and zero otherwise.
Table 7 reports the placebo regression results for the effect of SHHK Stock Connect on share price premium. In contrast to the baseline estimates, the coefficient on $FakePolicy$ is positive but insignificant across all columns. This finding suggests that share price premium is not driven by common time trend, the SHHK stock policy does affect the premium. Next step, we conducted the placebo test regarding whether there is an efficiency dependent effects of the SHHK Stock Connect. Table 8 reports a placebo test in which the SHHK Stock Connect is assumed to have been implemented two years later, in November 2016. The key coefficient of interest is the interaction term between the placebo policy indicator and market efficiency, $FakePolicy\times Effboth$.
The interaction term $FakePolicy\times Effboth$ is statistically insignificant, indicating that when the policy timing is artificially shifted, market efficiency does not systematically condition the relationship between the placebo policy and the AH price premium. This result suggests that the efficiency-dependent effect documented in the previous analysis is not driven by common time trends or spurious correlations, but is instead specific to the actual implementation of the SHHK Stock Connect.
Table 7: Placebo
@@ 4 Conclusion and Limitation
This study examines how the Shanghai–Hong Kong Stock Connect affects the A–H share price premium and whether the policy impact varies with market efficiency. Our findings are as follows: Firstly, the implementation of the Shanghai–Hong Kong Stock Connect significantly increases the AH premium by 18.4% on average. Second, the effect is heterogeneous and depends on the overall market efficiency. When we interact the policy indicator with market-efficiency measure, the marginal policy effect becomes pronounced for firms operating under lower market efficiency conditions, which increases the AH premium; on the other hand, the SHHK Stock Connect can reduce the AH premium for the dual-listing firms operating in higher market efficiency. Finally, we find no evidence of a premium response at the announcement stage. When April 2014 is used as an announcement event date, the announcement indicator and its interaction with efficiency are statistically insignificant, suggesting that the AH premium is unlikely to change prior to the implementation of the SHHK Stock Connect. This study also has two limitations. First, our sample is restricted to A–H firms whose A-shares are listed in Shanghai, so the evidence should be interpreted as specific to SHHK-eligible firms and may not generalise to Shenzhen-listed dual-listed firms. So, in the future, we may expand the sample to incorporate Shenzhen-listed A–H firms and examine whether the Shenzhen–Hong Kong Stock Connect exhibits a similar pattern. Second, while we find the Shanghai Hong Kong Stock Connect has different effects across firms, the mechanism explaining why the premium increases remains underdeveloped. Future research could introduce more direct channel variables, such as northbound and southbound holdings shares.

### v2

@@ 3 Empirical results
3.1 Baseline Effect of SHHK Stock Connect
We first examines the effect of the SHHK Stock Connect on the A-H share price premium. Table 2 reports the baseline dynamic panel regression results. Column (1) only includes the lagged premium and the SHHK Stock Connect indicator. The coefficient on SHHKPolicy is positive and statistically significant at the 1% level, indicating that the introduction of the Stock Connect is associated with an increase in the A-H share price premium.
Columns (2) to (6) progressively incorporate additional firm-level control variables, with Column (6) representing the full specification. Across all specifications, the estimated coefficient on SHHKPolicy remains positive and significant at 1% level. In the full model, the coefficient on SHHKPolicy is 0.184, implying that, holding other variables constant, the implementation of the SHHK Stock Connect is associated with an average increase of approximately 18.4% in the A-H share price premium for a given pair of dual-listed shares. The baseline regression results present an average effect of the SHHK Stock Connect on the A-H share price premium. However, the economic mechanisms under this effect remain unclear. If SHHK Stock Connect operates by alleviating market frictions
@@ 3.1 Baseline Effect of SHHK Stock Connect
and improving information transmission, its impact should be more pronounced in less efficient markets. This motivates our analysis of efficiency-dependent effects in the next subsection.
3.2 Market Efficiency-Dependent Effects of the SHHK Stock Connect
@@
Meng et al., (2023) finds that the SHHK Stock Connect increases the A-share market efficiency of AH dual listing firms through stock price informational efficiency and corporate governance. Motivated by their research, we test whether the effect of the SHHK Stock Connect on the A-H share price premium depends on market efficiency. We construct the following model:
$\text{Prem}_{i,t}=\alpha_{i}+\beta_{1}\text{Prem}_{i,t-1}+\beta_{2}\text{SHHKPolicy}_{t}+\beta_{3}\text{Effboth}_{t}+\beta_{4}\text{SHHKPolicy}\times\text{Effboth}_{t}+\gamma^{\prime}X_{it}+\epsilon_{it},$
where the $\text{Prem}_{i,t}$ is the A-H share price premium for share $i$ at month $t$. Follow Domowitz et al., (1997) and Fan and Wang, (2017), we control the lagged term $\text{Prem}_{i,t-1}$ to reduce the effect of autocorrelation.
[…]
@@
          Lagged Premium           0.633∗∗∗
          SHHKPolicy           -0.490∗∗
          Effboth           39.554∗∗
          SHHKPolicy $\times$ Effboth           -45.900∗∗∗
          Turnover           0.022∗∗∗
@@
          Demand           -0.009
          Firm Size           -0.108∗∗∗
          Interest Rate           0.007∗∗∗
@@
isolation. In this model, the marginal effect of SHHK Stock Connect is:
$\frac{\partial\text{Prem}}{\partial\text{SHHKPolicy}}=\beta_{2}+\beta_{4}\text{Effboth},$
so that, under our scaling where $\texttt{Effboth}<0$ and values closer to zero indicate higher efficiency, the estimated policy effect increases as Effboth becomes more negative (i.e., as efficiency declines). Consequently, the Stock Connect is associated with a larger change in the AH premium for firms operating under less efficient market conditions than for those with higher efficiency. Overall, the results in Table 3 support an efficiency-dependent policy effect both statistically and economically.
@@ 3.3 Robustness check
Table 5 examines whether the announcement of the SHHK Stock Connect has an efficiency-dependent effect on the A-H price premium. The coefficient on Announcement is negative but statistically insignificant (0.217-0.217), suggesting no significant change in the AH premium around the announcement date. More importantly, the interaction term Announcement $\times$ Effboth is also statistically insignificant (14.471-14.471), providing no evidence that the announcement effect varies systematically with market efficiency. Consistent with these results, the main effect of Effboth is not statistically different from zero in this specification. Overall, Table 5 indicates that the announcement itself and the interaction with market efficiency are unlikely to generate a measurable impact on the AH premium.
According to Meng et al., (2023), we use an alternative market efficiency proxy, Effboth2, as a robustness check. Specifically, for each trading day $t$ we compute the relative high-low spread as
$\text{S}_{2}=\frac{\text{H}_{t}-\text{L}_{t}}{\big(\text{H}_{t}-\text{L}_{t}\big)/2},$ (4)
@@
$\text{Eff}_{2}=-\text{S}_{2},$ (5)
For A-share market and H-share market, we calculate the $\texttt{Eff}_{2}$ separately and then add them up as an alternative proxy for overall market efficiency (EFFboth2). Table 6 presents our regression results for third robustness check. Even though SHHKPolicy is not statistically significant, the sign of coefficient is negative. The interaction term $\texttt{SHHKPolicy}\times\texttt{Effboth2}$ is negatively significant at 5% level. So, in general, this result is consistent with the result shown in Table 3. The SHHK Stock Connect was launched in November, 2014, to further test whether the effect comes from SHHK Stock Connect and assess the robustness of our previous findings, we conduct a placebo test by assigning the implementation of the SHHK Stock Connect to a later date. Specifically, we assume that the policy was introduced two years after its actual launch and construct a binary indicator, FakePolicy, which equals one
@@
for observations after November 2016 and zero otherwise.
Table 7 reports the placebo regression results for the effect of SHHK Stock Connect on share price premium. In contrast to the baseline estimates, the coefficient on FakePolicy is positive but insignificant across all columns. This finding suggests that share price premium is not driven by common time trend, the SHHK stock policy does affect the premium. Next step, we conducted the placebo test regarding whether there is an efficiency dependent effects of the SHHK Stock Connect. Table 8 reports a placebo test in which the SHHK Stock Connect is assumed to have been implemented two years later, in November 2016. The key coefficient of interest is the interaction term between the placebo policy indicator and market efficiency, $\texttt{FakePolicy}\times\texttt{Effboth}$.
The interaction term $\texttt{FakePolicy}\times\texttt{Effboth}$ is statistically insignificant, indicating that when the policy timing is artificially shifted, market efficiency does not systematically condition the relationship between the placebo policy and the AH price premium. This result suggests that the efficiency-dependent effect documented in the previous analysis is not driven by common time trends or spurious correlations, but is instead specific to the actual implementation of the SHHK Stock Connect.
Table 7: Placebo test
@@ 4 Conclusion and Limitation
This study examines how the Shanghai–Hong Kong Stock Connect affects the A–H share price premium and whether the policy impact varies with market efficiency. Our findings are as follows: Firstly, the implementation of the Shanghai–Hong Kong Stock Connect significantly increases the A-H premium by 18.4% on average. Second, the effect is heterogeneous and depends on the overall market efficiency. When we interact the policy indicator with market-efficiency measure, the marginal policy effect becomes pronounced for firms operating under lower market efficiency conditions, which increases the A-H premium; on the other hand, the SHHK Stock Connect can reduce the A-H premium for the dual-listing firms operating in higher market efficiency. Finally, we find no evidence of a premium response at the announcement stage. When April 2014 is used as an announcement event date, the announcement indicator and its interaction with efficiency are statistically insignificant, suggesting that the A-H premium is unlikely to change prior to the implementation of the SHHK Stock Connect. This study also has two limitations. First, our sample is restricted to A–H firms whose A-shares are listed in Shanghai, so the evidence should be interpreted as specific to SHHK-eligible firms and may not generalise to Shenzhen-listed dual-listed firms. So, in the future, we may expand the sample to incorporate Shenzhen-listed A–H firms and examine whether the Shenzhen–Hong Kong Stock Connect exhibits a similar pattern. Second, while we find the Shanghai Hong Kong Stock Connect has different effects across firms, the mechanism explaining why the premium increases remains underdeveloped. Future research could introduce more direct channel variables, such as northbound and southbound holdings shares.